OpenStudy (anonymous):

5(x − 1) = 3(x + 2) − 7

OpenStudy (anonymous):

Solve for x?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Well first we have to distribute out what is in the parentheses. Can you do that? What are you left with?

OpenStudy (anonymous):

5x-5= 3x+6-7 right ?

OpenStudy (anonymous):

Yes good job!

OpenStudy (anonymous):

the right hand side can be simplified from\[5x-5= 3x+6-7 \]to \[5x-5= 3x-1 \]

OpenStudy (anonymous):

now we try to get all the x to one side and all the other numbers on the other sides let's keeps x on the left side so let's get rid of the -5 on the left side by adding 5 to both sides

OpenStudy (anonymous):

\[5x−5=3x−1\]\[5x−5+5=3x−1+5\]\[5x=3x+4\]subtract -3x to both sides to rid of the 3x on the right hand side. \[5x=3x+4\] \[5x-3x=3x-3x+4\]\[2x=4\]

OpenStudy (anonymous):

From where I left of, divide both sides by 2 and you get your answer

OpenStudy (anonymous):

What is x?

OpenStudy (anonymous):

thank you so much !

OpenStudy (anonymous):

Yup np